xx NullDecomp

NullDecomp

Yosef Zlochower <yosef@phys.utb.edu>

May 23 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11storexylms 10Scope: privatex01 BOOLEAN

11Description: should the Ylm’s be stored rather than recomputed each time?

11              10  x01  Default: yes

11

11usexrsylm  10Scope: privatex01  BOOLEAN

11Description: use the rsYlm’s rather than the standard sYlm’s

11              10  x01  Default: no

11

11lxmax  10Scope: restrictedx01  INT

11Description: Decompose into spherical harmonics \ell = spin up to \ell = lxmax

11 Range 10 x01 Default: 9
112:  102 or greater

11

11nxangxevxoutsidexeq 10Scope: shared from NULLGRIDx01INT

11

11nxangxghostxpts10Scope: shared from NULLGRIDx01INT

11

11nxangxptsxinsidexeq 10Scope: shared from NULLGRIDx01INT

11

11nxangxstencilxsize 10Scope: shared from NULLGRIDx01INT

11

7 Interfaces

General Implements: nulldecomp

Inherits: nullgrid Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 kern 10 kern x01 compact110
11                 10  x01 description11kernel of the integral
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 area 10 area x01 compact110
11                 10  x01 description11integral algorithms internal variable
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullDecomp. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:11x
11kern area 11x

11x  11x

Scheduled Functions

CCTKxBASEGRID x nulldecompxstartup x initialization x

11x11After: 11nullgridxsetup
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function